# Applications of Generative AI (GAI) for Mobile and Wireless Networking: A Survey

### v1

@@ IV-E Lessons Learned and Keys Takeaway
In addition, recent developments in GAI have proven crucial in the fight against jamming attempts that interfere with wireless communication. GAI improves malware detection and stops the propagation of threats by analyzing network data and device behavior to quickly identify and isolate infected devices. Also, by encrypting data and disguising it within seemingly innocent material, steganography’s novel application of GAI enables secure data transmission. Concurrently, biometric-based authentication that uses GAI strengthens identity verification processes by recognizing distinctive biometric patterns like fingerprints or facial features. These uses serve as a reminder of the important of GAI for improving data security and facilitating safe and secure user verification.
Mobile networking offers more robust security measures due to the adaptability of GAI in varied security concerns. GAI approaches hold promise to increase the state-of-the-art in mobile network security and communication and data transmission integrity by strengthening defenses against sophisticated attacks. In addition, our review uncovered numerous innovative research methods that further demonstrate the potential of GAI to offer a thorough defence against changing threats and improve the mobile networking security environment, as summarized in Table IV. The future of mobile networking security is bright due to ongoing developments in GAI, which will ensure a safer and more secure digital environment.
In summary, the key takeaways from this section are as follows.
  • GAI strengthens wireless security against anomalies and intrusions in real time.
  • GAI combats jamming attacks and ensures secure data transmission with privacy protection.
  • Biometric-based authentication and malware detection are reinforced using GAI techniques.
  • Ongoing research in this area provides more resilient security measures for mobile networking.
@@ V Application of GAI in Semantic Communication
SemCom revolutionizes communication by addressing limitations faced by conventional systems, even with advanced techniques. It considers the semantics of transmitted data, allowing imperfect decoding while fulfilling the intended purpose and reducing unnecessary data traffic and energy consumption. SemCom presents three primary challenges: exploiting shared knowledge, efficient encoding and decoding, and data aggregation. Overcoming these challenges will be a breakthrough step in offering new alternatives in representation learning and disentanglement learning, accelerating the realization of AI-driven tasks and recommendation systems. Fig. 6 illustrates the general application of GAI in SemCom.
Figure 6: Illustration of GAI in semantic communication. The GAI take place in three blocks, i.e., Knowledge Extractor, DJSCC Encoder and Decoder. The Knowledge Extractor aims to generate both common knowledge, and variant data from the original data. The DJSCC aim to compress and reconstruct the data with a high compression ratio and robust to noise.
@@ V-A Knowledge Abstraction
Knowledge abstraction is crucial for representing abstract knowledge among users and acquiring shared knowledge from distributed data across device domains. KG are the most prominent candidate among contemporary techniques for this purpose. However, manually constructing KGs is labor-intensive and impractical for scalability [130, 131]. Toward automatic KG construction, three FL-enabled designs for GAI deployment in wireless networks have been recently proposed [43]. The success of these design lines is discussed using three key knowledge techniques.
@@ V-A1 Knowledge Summarization
Summarization techniques aim to create concise abstracts that retain essential information for logical reasoning, omitting less crucial details. Therefore, leveraging Language Models (LMs) to extract knowledge from KGs has recently been prevalent. However, automatically learning optimal prompts typically requires abundant training data [132, 133, 134], which may not have useful meaning to new relationships. Despite its success with GPT-3, the approach in [135] is not transferable to other LMs due to its reliance on ad hoc bounded learning and the scalability of GPT-3.
To address the limitations of the aforementioned methods, BertNet [136] proposes an unsupervised approach to automatically generate diverse alternate prompts while requiring the input of minimal relationship definitions. Compared to conventional approaches that rely on extensive human-annotated data or large KGs, BertNet enables extraction of knowledge about new relationships previously inaccessible.
@@ V-A2 Knowledge Representation Optimization
This approach aims to enhance knowledge representation efficiency by transforming complex rule-based structures into more compact formats, such as ontologies or Binary Decision Diagrams (BDDs). A previous study  [137] introduces a spatial-functional embedding for KG completion that represents entity pairs as points and relations as hyper-parallelograms in $\mathbb{R}^{2d}$. This enables pattern representation through the spatial relationship of hyper-parallelograms, providing a clear geometric interpretation of embeddings and captured patterns.
Despite extensive research on knowledge extraction from data, creating representations for relation extraction in a KG remains challenging, especially with distantly labeled datasets due to significant label noise. To tackle this circumvent, a solution called LERP was proposed [138]. LERP is an efficient framework capable of learning prototypes for each relation based on contextual information, capturing the intrinsic semantics of these relations by observing noisy labels.
On the other hand, summarizing KGs encounter two critical problems: limited observations for model training and evolving characteristics of novel entities in temporal KGs. To tackle mose issues, [139] presents MetaTKGR, which employs meta-learning to adjust domain adaptation over time.
@@ V-A3 Knowledge Aggregation
Aggregation techniques combine similar or related knowledge from diverse data sets or learned models to reduce redundancy and strengthen decision-making or the ability to adapt more effectively to new domains [140]. Nevertheless, exploiting automatic knowledge abstraction via entity retrieval encounters three primary obstacles: 1) Context and entity affinity primarily rely on vector dot products, possibly overlooking fine-grained interactions. 2) Large entity sets require a significant memory footprint to store dense representations. 3) Training necessitates subsampling a suitably challenging set of negative data.
Driven by these challenges, a solution called Generative Entity Retrieval (GENRE) is introduced [141] to use a transformer-based generative model in conjunction with a pre-trained NLP model. The unique feature of GENRE lies in its implementation of a constrained decoding strategy, ensuring that every generated entity belongs to a predefined candidate set. As a consequence, the outputs of the GAI model are less likely to deviate from the desired entities.
@@ V-B1 Structured and Compact Data Representation
Loss Function Design. This effort focuses on formulating an efficient loss function to minimize the discrepancies between the original and reconstructed data while capturing noise features from the wireless channel. However, revisiting the classical Shanon theorem [144] indicates that, for a memoryless communication channel, optimal communication involves two distinct processes as the data length approaches infinity:
@@
Therefore, recent efforts focus on leveraging DL and GAI to JSCC system with acceptable approximation outcomes. Neural Error Correcting and Source Trimming (NECST) is the first effort in this cutting-edge area [145]. This advanced DL framework compresses and corrects input images within a set bit-length constraint. NECST surpasses traditional models by integrating a discrete channel simulation, injecting noise directly into latent representations to enhance resilience in challenging channel conditions. When applied to the Deep JSCC (DJSCC) system [146], NECST activates self-learning, swiftly adapting to new environments without significant computational overheads from training anew. The work in [147] tackles the challenge of JSCC over a discrete-input Additive White Gaussian Noise (AWGN) channel by mapping input images to codewords and decoders to minimize average end-to-end distortion. The specific problem of the end-to-end JSCC is regarded by a quantizer in the middle, where features are extracted from input images and quantized to constellation points through a DNN-based encoder. After transmission, the receiver reconstructs the image from noisy observations using another DNN. The study [148] focuses on mitigating data noise from an impaired channel by capturing noise statistics and applying noise reduction based on channel behavior. Meanwhile, the work in [149] introduces a GAN-based approach for improved data reconstruction in lossy compression. Through a regularization function, this method can address both universal rate distribution and perception functions concurrently.
@@
With the demands of short training time and lower computational power, achieving high compression ratios without compromising data reconstruction, especially in 3D-based AI tasks like point cloud representations, is not a straightforward task. Addressing this, DeepPCAC, a novel approach that directly encodes and decodes point cloud attributes, has been introduced [153]. This method surpasses traditional compressed methods due to adopting geometry instead of using voxelization or point projection.
As previously mentioned, the advent of GOSC signifies the onset of a new era in lossy compression within the field of SemCom, as it does not require reconstructing the entire dataset. Instead, it compresses incoming data as long as AI task performance is achieved. This approach is effective as the AI era enters a prosperous phase, with machine-to-machine and human-to-machine communication becoming more prevalent in IoT Networks [28]. The main feature of GOSC is to extract invariant features among domains [154] or classes [155]. Therefore, as long as the encoders can capture invariant features, they identify crucial AI task features. Based on foundational principles, the authors in [156] integrate invariant learning in a lossy compressor, preserving lossless prediction in AI classification. This method encapsulates essential features for AI tasks, enabling lossy compression while maintaining consistent AI performance.
Architecture Design. The main focus of this effort shifts to the design of an efficient model architecture aimed at enhancing data reconstruction efficiency irrespective of the applied loss function. A typical representative of capturing essential data representations in various media formats is the Attention mechanism, known as Transformer [61]. On this basis, Deep SemCom (DeepSC) architecture for speech recognition applications has been developed to address noise and distortion challenges through joint semantic-channel coding [157]. Similarly, the upgrade versions, U-DeepSC and MUDeepSC [158, 159], accommodate text and image transmission and facilitate semantic information extraction. In that, a layer-wise knowledge transfer technique is exploited to enhance information exchange across tasks, similar to the UNet architecture [160, 161, 162], ensuring detailed information preservation.
Another approach is to design an efficient model architecture for data compression and reconstruction by leveraging the robustness of model diffusion. According to [163], diffusion models comprise three primary components the forward process, the reverse process, and the sampling procedure, which affect the probability distribution of a specified dataset. These models deduce the latent structure of a dataset by capturing the way in which data points propagate through their latent space [164]. Upon the conceptualization of the SemCom process as a sequential Gaussian model [165], the work in [166] introduces an improved lossy compression method capable of seamlessly integrating diffusion models. This method utilizes noise from the SemCom process’s communication channel to improve compression and semantic encoding, enhancing both communication efficiency and data reconstruction noticeably.
@@ V-B2 Knowledge-assisted Compression
Knowledge-Assisted Compression (KAC) is a recent groundbreaking compression approach that inherits from knowledge-based reasoning, as mentioned in Section V-A3. Knowledge-assisted compression approaches tackle the process of compressing or compacting knowledge representations to improve reasoning efficiency and storage requirements. In KAC, systems use explicit knowledge, often represented as rules, facts, or ontologies, to conclude, make inferences, or solve problems. Thus, with the aid of knowledge representation, the compressor and reconstructor can further improve the compression efficiency with the common knowledge among data.
In [167], the authors propose the Residual Compressor (RC) method, akin to knowledge-aided SemCom, where the BPG lossy compression algorithm is used to generate a versatile representation set. This enables subsequent lossless compression and discovers fixed lost information. In addition, a new network architecture with two-stage execution is provided to approximate lost information from lossy compression, which is not only beneficial for reconstructing compressed data but also retains all information. The authors in [168] leverage the generative AI model to transform the original data $S$ into another form of data $\widetilde{S}$. In other words, the generative model $f_{K}(\cdot)$ aggregates the source data with the background knowledge, known by both the transmitters and receivers.
Analogical reasoning is a robust alternative to make reasonable use of KGs in decoding data representation. Fundamental, analogical reasoning aims at increasing machine intelligence by associating data representation (e.g., CV, NLP) with relational, structural, and analogical reasoning. However, research on knowledge-assisted data reconstruction seems to have little focus on knowledge structure. On that basis, a new method for building KGs has been proposed in [169], where Graph Neural Networks (GNN) are considered to develop an efficient data reconstruction technique. Reconstructing data from a single task’s knowledge and data representation may not apply to compressed data operations from multiple semantic knowledge sources. To address this, the work in [170] introduces a competitive embedding framework for logical reasoning over KGs. This approach leverages structure mapping theory and a Transformer-based GAI model to aggregate multimodal knowledge for link prediction, enabling seamless integration into the data reconstruction process. The work in [171] discovers two types of information redundancy in remote sensing image compression: pixel-wise and non-Region-Of-Interest (nROI) data. This allows a more favorable bit allocation strategy by assigning a greater number of bits to the ROI while allocating fewer bits to the nROI. Likewise, the authors in [172] introduce a task-oriented image compression framework that can extract and process ROI and nROI data independently. This framework significantly enhances accuracy and compression efficiency.
@@
To address the oversight of semantic noise and system robustness in data encoding and compression, two efficient methods of Masked Autoencoder (MAE) [173] and Vector Quantized VAE (VQ-VAE) [174] are introduced. Since these methods utilize the GAI model to generate noise and pre-learning adversarial risks before transmission, they not only demonstrate robustness against semantic noise, but also significantly enhance semantic reconstruction via knowledge-based encoder-decoder architectures compared to conventional techniques. The work in [175] presents a quantization-based knowledge extractor, where an automatic mask generation based on semantic knowledge information is developed to regulate the compression ratio of transmitted data. By doing so, the approach enables the SemCom system to dynamically adapt to variable-length coding systems. To further improve compression efficiency within the SemCom system introduced in [157], the authors of [176] employ LSTM techniques to formulate a memory-aided SemCom system. With historical information about the data and compression efficiency, this approach yields higher compression efficiency without compromising communication efficiency, meaning there is no need to transmit additional data via the communication channel.
@@
The research in [177, 178] introduces a new SemCom framework aimed at reliable communication with minimal data transfer across diverse services and users. Central to this framework are GFlowNets, inspired by GAI networks, which enable causal reasoning for SemCom. These networks serve as causal describers to highlight transitions in task distributions with changing contexts when using GFlowNets to identify task event transitions and form a graph for GNN to improve data reconstruction. The study in [179] introduces a causal semantic communication method for abstracting knowledge to enhance communication efficiency among distributed clients. In that, a generative AI model is deployed with the ability to produce causally invariant representations, facilitating the SemCom system to prioritize transmission of the most meaningful features to the goal-oriented task at the recipient’s end. The research in [180] proposes a knowledge-enhanced SemCom framework in which the receiver can more actively utilize the data in the knowledge base for semantic reasoning and decoding, based on only affecting the parameters rather than the structure of the neural networks at the transmitter. This is achieved using a transformer-based knowledge extractor to find relevant factual triples for the received noisy signal.
V-C Generative AI and Distributed Learning
GAI has demonstrated remarkable robustness in enhancing performance in distributed learning. To be specific, GAI techniques produce generative data representations that extract generalized knowledge or meaningful features, supporting training on distributed local devices without requiring access to unseen data domains.

### v2

@@ IV-E Lessons Learned and Key Takeaways
In addition, recent developments in GAI have proven crucial in the fight against jamming attempts that interfere with wireless communication. GAI improves malware detection and stops the propagation of threats by analyzing network data and device behavior to quickly identify and isolate infected devices. Also, by encrypting data and disguising it within seemingly innocent material, steganography’s novel application of GAI enables secure data transmission. These uses serve as a reminder of the importance of GAI for improving data security and facilitating safe and secure user verification.
Mobile networking offers more robust security measures due to the adaptability of GAI in varied security concerns. GAI approaches hold promise to increase the state-of-the-art in mobile network security and communication and data transmission integrity by strengthening defenses against sophisticated attacks. Various use cases of GAI approaches in wireless network security can be compiled in Table IV. It demonstrates how several GAI models solve particular security issues, improving network resilience, threat detection, and data protection. In addition, our review uncovered numerous innovative research methods that further demonstrate the potential of GAI to offer a thorough defence against changing threats and improve the mobile networking security environment, as summarized in Table V. The future of mobile networking security is bright due to ongoing developments in GAI, which will ensure a safer and more secure digital environment.
The key takeaways from this section are as follows.
  • Key 1. GAI strengthens wireless security against anomalies and intrusions in realtime.
  • Key 2. GAI combats jamming attacks and ensures secure data transmission with privacy protection.
  • Key 3. GAI achieves enhanced information encryption through the steganography technique.
  • Key 4. Ongoing research in GAI provides more resilient security measures for mobile networking.
@@ V Application of GAI in Semantic Communication
SemCom revolutionizes communication by considering the semantics of transmitted data to allow imperfect decoding while fulfilling the intended purpose and reducing unnecessary data traffic and energy consumption. However, SemCom also imposes three challenges: 1) generating common knowledge, 2) exploiting shared knowledge for efficient encoding and decoding, and 3) facilitating communication-free knowledge sharing among clients in distributed learning. Following that, we first discuss how to generate common knowledge in SemCom. Next, we describe the strategies for integrating this common knowledge into data compression and reconstruction design. Afterward, we explore the role of GAI in accelerating the execution of AI-driven tasks and recommender systems in distributed networking, such as a knowledge generator and domain-invariant representation learning across devices. Fig. 7 illustrates a basic block of GAI in SemCom. Finally, we conclude this section with useful lessons and key takeaways.
Figure 7: Illustration of GAI in semantic communication. The GAI take place in three blocks, i.e., Knowledge Extractor, DJSCC Encoder and Decoder. The Knowledge Extractor aims to generate both common knowledge and variant data from the original data. The DJSCC aim to compress and reconstruct the data with a high compression ratio and robust to noise.
@@ V-A Knowledge Abstraction
Knowledge abstraction is a technique used to represent abstract knowledge among users and acquire shared knowledge from distributed data across device domains. However, manually constructing KGs is labor-intensive and impractical for scalability [143, 144]. Toward automatic KG construction, three FL-enabled GAI designs in wireless networks have been recently proposed [40], with three key knowledge techniques.
@@ V-A1 Knowledge Summarization
Summarization techniques create concise abstracts that retain essential information for logical reasoning, omitting less crucial details. This results in the prevalence of leveraging Language Models (LMs) to extract knowledge from KGs. However, automatically learning optimal prompts typically requires abundant training data [145, 146, 147], which may not have useful meaning to new relationships. Despite its success with GPT-3, the approach in [148] is not transferable to other LMs due to its reliance on ad hoc bounded learning and the scalability of GPT-3.
To address the limitations of the aforementioned methods, BertNet [149] proposes an unsupervised learning approach to automatically generate diverse alternate prompts while requiring the input of minimal relationship definitions. Compared to conventional approaches that rely on extensive human-annotated data or large KGs, BertNet enables extraction of knowledge about new relationships previously inaccessible.
@@ V-A2 Knowledge Representation Optimization
This approach enhances knowledge representation efficiency by transforming complex rule-based structures into more compact formats, such as ontologies or Binary Decision Diagrams (BDDs). A previous study  [150] introduces a spatial-functional embedding for KG completion that represents entity pairs as points and relations as hyper-parallelograms in $\mathbb{R}^{2d}$. This enables pattern representation through the spatial relationship of hyper-parallelograms, providing a clear geometric interpretation of embeddings and captured patterns.
However, creating representations for relation extraction in a KG remains challenging due to significant label noise in distantly labeled datasets. To tackle this circumvent, a solution called LERP was proposed [151]. LERP is an efficient framework capable of learning prototypes for each relation based on contextual information, capturing the intrinsic semantics of these relations by observing noisy labels. On the other hand, summarizing KGs also encounter two critical problems: limited observations for model training and evolving characteristics of novel entities in temporal KGs. This motivates the presence of MetaTKGR in [152], which employs meta-learning to adjust domain adaptation over time.
@@ V-A3 Knowledge Aggregation
This technique refers to the combination of similar and related knowledge from diverse data sets or learned models to reduce redundancy and enhance decision-making or the adaptive ability to new domains [153]. Nevertheless, exploiting automatic knowledge abstraction via entity retrieval encounters three primary obstacles: 1) Context and entity affinity primarily rely on vector dot products, possibly overlooking fine-grained interactions; 2) Large entity sets require a significant memory footprint to store dense representations; and 3) Training necessitates subsampling a suitably challenging set of negative data.
Driven by these challenges, a solution called Generative ENtity REtrieval (GENRE) is introduced [154], which combines a transformer-based generative model and a pre-trained NLP model. GENRE offers implementations with constrained decoding, ensuring that every generated entity belongs to a predefined candidate set. As a consequence, the outputs of the GAI model are less likely to deviate from the desired entities.
@@ V-B1 Structured and Compact Data Representation
Loss Function Design. This effort focuses on formulating an efficient loss function to minimize discrepancies between the original and reconstructed data while capturing noise features from the wireless channel. However, revisiting the classical Shanon theorem [157] indicates that, for a memoryless communication channel, optimal communication involves two distinct processes as the data length approaches infinity:
@@
Therefore, recent efforts focus on leveraging DL and GAI to JSCC system with acceptable approximation outcomes. Neural Error Correcting and Source Trimming (NECST) is the first effort in this cutting-edge area [158]. This advanced DL framework compresses and corrects input images within a set bit-length constraint. NECST surpasses traditional models by integrating a discrete channel simulation, injecting noise directly into latent representations to enhance resilience in challenging channel conditions. When applied to the Deep JSCC (DJSCC) system [159], NECST activates self-learning, swiftly adapting to new environments without significant computational overheads from training anew. The work in [160] tackles the challenge of JSCC over a discrete-input Additive White Gaussian Noise (AWGN) channel by mapping input images to codewords and decoders to minimize average end-to-end distortion. The specific problem of the end-to-end JSCC is regarded by a quantizer in the middle, where features are extracted from input images and quantized to constellation points through a DNN-based encoder. After transmission, the receiver reconstructs the image from noisy observations using another DNN. The study [131] focuses on mitigating data noise from an impaired channel by capturing noise statistics and applying noise reduction based on channel behavior. The work in [161] introduces a GAN-based approach for improved data reconstruction in lossy compression. Through a regularization function, this method can address both universal rate distribution and perception functions concurrently.
@@
With the demands of short training time and lower computational power, achieving high compression ratios without compromising data reconstruction, especially in 3D-based AI tasks like point cloud representations, is not a straightforward task. Inspired by this, DeepPCAC, a novel approach that directly encodes and decodes point cloud attributes, is introduced [165] to surpass traditional compressed methods due to adopting geometry instead of using voxelization or point projection.
As mentioned earlier, the advent of GOSC signifies a new era in lossy compression within SemCom due to no need for reconstructing the entire dataset. Instead, it compresses incoming data as long as AI task performance is achieved. This approach is effective as the AI era enters a prosperous phase, with machine-to-machine and human-to-machine communication becoming more prevalent in IoT Networks [44]. The main feature of GOSC is to extract invariant features among domains [166] or classes [167]. Therefore, as long as the encoders can capture invariant features, they identify crucial AI task features. Based on foundational principles, the authors in [168] integrate invariant learning in a lossy compressor, preserving lossless prediction in AI classification. This method encapsulates essential features for AI tasks, enabling lossy compression while maintaining consistent AI performance.
Architecture Design. There are two directions for this regard. The first focus is to design an efficient model architecture to enhance data reconstruction efficiency. Based on the Attention mechanism [53], Deep SemCom (DeepSC) architecture for speech recognition applications is able to address noise and distortion challenges through joint semantic-channel coding [169]. Similarly, the upgrade versions, U-DeepSC and MUDeepSC [170, 171], accommodate text and image transmission and facilitate semantic information extraction. In that, exploiting a layer-wise knowledge transfer technique enhances information exchange across tasks, similar to UNet architectures [172, 173, 174], ensuring information preservation.
The second focus is to design an efficient model architecture for data compression and reconstruction. The use of diffusion models [175] helps deduce the latent structure of a dataset by capturing how data points propagate through their latent space [176]. Upon conceptualizing the SemCom process as a sequential Gaussian model [177], the work in [178] introduces an improved lossy compression method capable of seamlessly integrating diffusion models. This method utilizes noise from the SemCom process’s communication channel to improve compression and semantic encoding, enhancing communication efficiency and data reconstruction noticeably.
@@ V-B2 Knowledge-assisted Compression
Knowledge-Assisted Compression (KAC) is a recent groundbreaking compression approach that inherits from knowledge-based reasoning, as mentioned in Section V-A3. Knowledge-assisted compression approaches tackle the process of compressing or compacting knowledge representations to improve reasoning efficiency and storage requirements. KAC systems use explicit knowledge, often represented as rules, facts, or ontologies, to conclude, make inferences, or solve problems. Thus, the compressor and reconstructor can further improve the compression efficiency with the common knowledge among data.
@@
Noise and Channel Knowledge. To address the oversight of semantic noise and system robustness in data encoding and compression, two efficient methods of Masked Autoencoder (MAE) [179] and Vector Quantized VAE (VQ-VAE) [180] are introduced, where GAI is exploited to generate noise and pre-learning adversarial risks before transmission. Thus, both achieve robustness against semantic noise and enhanced semantic reconstruction via knowledge-based encoder-decoder architectures simultaneously. The work in [181] presents a quantization-based knowledge extractor, where an automatic mask generation based on semantic knowledge information is developed to regulate the compression ratio of transmitted data. By doing so, the approach enables the SemCom system to dynamically adapt to variable-length coding systems. To further improve compression efficiency within the SemCom system introduced in [169], the authors of [182] employ LSTM techniques to formulate a memory-aided SemCom system. With historical information about the data and compression efficiency, this approach yields higher compression efficiency without compromising communication efficiency, meaning there is no need to transmit additional data via the communication channel. Another approach is to leverage the noise and channel feedback from the receivers to perform estimation of the ideal data transmission at the encoder to mitigate the channel noise [183, 184, 185, 186, 187, 188].
Human-defined Knowledge. The authors in [189] introduced one of the earliest approaches to incorporating language-based knowledge into data compression. This study proposes a framework wherein a predefined library of language knowledge, such as an electronic dictionary, is accessible. The decompression process can then selectively utilize relevant knowledge from this library to enhance compression efficiency significantly. Thus, this approach not only underscores the potential of leveraging language-based knowledge for data reconstruction but also achieves a relatively simple design.
NLP-based Knowledge Generator. This approach can be divided into two folds. The first approach utilizes pretrained LLMs to generate auxiliary representations that enhance the performance of the decoder [190, 191], with two advantages. LLMs trained on extensive datasets produce representations that encapsulate a wealth of knowledge and no additional training is required, making the system easy to integrate. The second approach is to leverage prompting from the model of BLIP and CLIP to improve the decoder performance [192, 193]. This facilitates huge computation performance for inference time without overparameterization typical in LLMs.
@@
Causality-Invariant Representation. This approach enables practical causal reasoning in SemCom, recently attracting considerable attention [194]. The generalized characteristics of causality invariance benefit radio nodes in learning and establishing a stable representation for a specific semantic content element [195, 196, 197]. Once this representation is formed, radio nodes can use it consistently to describe semantic content without being affected by variations in distribution, domain, or context. Given causality invariant representations, causality-based neural symbolic AI can indeed help achieve minimal representations, create symmetric communication channels [198], enable generalizability, and reduce data transmitted. For example, SemCom frameworks in [198, 199] can achieve reliable communication with minimal data transfer across diverse services and uses. With an emphasis on GFlowNets, such GAI-inspired networks serve as the basis means in highlighting task distributions with changing contexts to identify task event transitions and form a graph for GNN to improve data reconstruction. Similarly, using a causal SemCom method enhances communication efficiency among distributed clients significantly [200]. In that, GAI models with causally invariant representation abilities help the SemCom system prioritize the transmission of the most meaningful features to the goal-oriented task at the recipient’s end. Another example is employing a knowledge-enhanced SemCom framework [201] to help the receiver actively utilize the data in the knowledge base for semantic reasoning and decoding. This is achieved using a transformer-based knowledge extractor to find relevant factual triples for the received noisy signal.
V-C Generative AI in Distributed Learning
GAI has demonstrated remarkable robustness in enhancing performance in distributed learning. To be specific, GAI techniques produce generative data representations that extract generalized knowledge or meaningful features, supporting training on distributed local devices without requiring access to unseen data domains. In FedGen [202], the server based on local model performance cooperatively trains a GAI model as a lightweight generator to amass user information and that is capable of generating pseudo data representation without accessing their true data (see Fig. 9). Thus, each local user can have the additional knowledge of other distributed clients, without sharing data among distributed devices. This method is robust as it can significantly enhance the performance and communication efficiency of FL, all while ensuring that user data remain secure and private. In [41], an AaaS framework based on AIGC models is designed to address the challenges in training and deployment overheads for edge networks.